OpenStudy (anonymous):

A ladder is leaning against a wall. The ladder touches the ground 15 inches from the base of the wall is 25 degrees. The angle between the base of the wall and the ground is a right angle. What is the length of the ladder?

OpenStudy (anonymous):

Can anyone please help me( this lonely helpless girl)

OpenStudy (cwrw238):



OpenStudy (anonymous):

Isn't it cosine Sorry my post will be hard to put up because I'm typing from a phone

OpenStudy (cwrw238):

is that correct?

OpenStudy (cwrw238):

yes you use the cosine

OpenStudy (cwrw238):

cos 25 = 15 / x x = 15 / cos 25

OpenStudy (anonymous):

Okay okay I'm back now on the iPad u guys have to post up things one time at a time my connection is getting worse by the minute